OpenStudy (anonymous):

Solve for x: (x - 9)2 = 1 So far I got (x-9)2=1 (√x-92)=√1 x-9=1 x=1=9 x=10 was told to...take the square root of both sides. So x - 9 = +-1. Then what? I'm confused.

hero (hero):

Hint: \[(x-9)^2 = 1\]\[x - 9 = \pm \sqrt{1}\]

OpenStudy (anonymous):

x=10,8

hero (hero):

In general, when \[a^2 = b\], \[a = \pm \sqrt{b}\]

OpenStudy (p0sitr0n):

\[\sqrt{a^2} = \left| a \right|\]

OpenStudy (radar):

Expand getting: \[x ^{2}-18x + 81 = 1\]\[x ^{2}-18x + 80=0\]

OpenStudy (radar):

\[(x-8)(x-10)=0\]\[x=10\]\[x=8\]

OpenStudy (anonymous):

Thanks.